# Entanglement and apparent thermality in simulated black holes

### v1

@@ Abstract
We investigate the apparent thermality of Hawking radiation in the semi-classical limit of quantum black holes using the mean-field limit of a chiral spin-chain simulator, which models fermions propagating on a black hole space-time in the continuum. In this free-theory regime, no genuine thermalisation occurs. Nevertheless, we show that a bipartition across the event horizon yields a reduced density matrix whose mode occupations follow an apparent thermal Fermi–Dirac distribution. In contrast, partitions away from the horizon do not exhibit thermal behaviour, reflecting the absence of true equilibration. Our results demonstrate that Hawking radiation appears thermal only with respect to horizon bipartitions in free theories, while true thermal behaviour emerges only in the presence of interactions deep in the black hole interior.
@@ I Introduction
Despite theoretical advancements in the study of quantum black holes, there remains several fundamental open questions, including whether, via black hole formation and evaporation, information of the black hole’s internal content is erased (the black hole information paradox) [7, 8]; why black hole entropy appears to be independent of the number of quantum fields [9, 10, 11]; what quantum degrees of freedom can be attributed to black hole entropy [12, 13]. In an effort to illuminate our understanding of these questions, there have been significant efforts, theoretically and, more recently, experimentally [14, 15, 16], towards realising physical analogues of black holes. Such a system was first proposed by Unruh, who argued that sonic analogues of black holes would experience a process analogous to the Hawking effect [17, 18]. Since then, a myriad of systems have been proposed to simulate not only the physics of black holes but also cosmological inflation, the early universe, and aspects of the standard model [19, 20, 21, 22, 23].
@@
Here, we consider a chiral spin-chain model [24, 25, 26, 27] that, in its continuum limit, due to the emergence of an effective Lorentz invariance in its low-energy limit, describes a theory of Dirac fermions in a curved space-time geometry. This Lorentz invariance is naturally broken at the length scales of the spin-chain’s lattice spacing as a consequence of the model’s non-linear dispersion relation. We calculate the density of states of the fermion zero-modes in various sub-regions of the space-time describing a (1+1)-dimensional black hole, which is finite as a consequence of the non-linear dispersion relation [22, 23, 21, 28, 29, 30, 31, 32]. By assuming that the fermions inside the black hole are in thermal equilibrium at the Hawking temperature $T_{H}$, the density of states can be used to compute the entropy of the fermion zero modes across different regions of the black hole space-time. This includes the interior, where the entropy diverges logarithmically as the lattice spacing approaches zero [33, 34, 35], as well as the exterior.
@@
In parallel, we employ numerical simulations to investigate the entanglement entropy of the chiral spin-chain following the partitioning of the lattice into two subsystems. Generally, we expect that the entropy of the fermion zero-modes agrees with the entanglement entropy of the chiral spin-chain if the state describing the partitioned system is thermal. Indeed, in this case, the von Neumann entropy is equivalent to the thermodynamic entropy of the subsystem’s microstates [13, 36]. When the system is partitioned at the event horizon, we find that the entanglement entropy is well described by the equation for the fermion zero-mode entropy of the black hole’s interior, suggesting that the ground state of the partitioned system is thermal. However, when the system is partitioned at a lattice site further from the horizon, we find that this agreement breaks down, indicating that the subsystem’s ground state appears thermal only when partitioned near the event horizon. We verify this by studying the subsystem’s mode occupation, which reveals a thermal Fermi-Dirac distribution only when the system is partitioned at the horizon.
@@ II Lattice simulator of Dirac fermions in black hole background
In this section, we introduce the chiral spin-chain model [24, 25, 26, 27], for which, due to the emergence of an effective Lorentz invariance at low energies, the continuum limit Hamiltonian is described by that of massless Dirac fermions propagating on a $(1+1)$-dimensional curved space-time geometry. Thus, by simulating the chiral spin-chain Hamiltonian, we can probe the entropy and thermality of fermions on a black hole space-time. Whilst the derivation of this continuum limit is outlined in this section, it is detailed further in Appendix A..
@@
By re-labelling the lattice sites such that they alternate between two sub-lattices, $A$ and $B$, and introducing a two-site unit cell, the mean-field Hamiltonian of Eq. (2) can, following a Fourier transform, be expressed as [24, 25, 26, 27]
@@
$E(p)=v\sin(a_{c}p)\pm u\sqrt{2+2\cos(a_{c}p)},$ (5)
which has Fermi points located at momenta $p_{0}=\frac{\pi}{a_{c}}$ and $p_{\pm}=\pm\frac{1}{a_{c}}\arccos(1-\frac{2u^{2}}{v^{2}})$, of which only the former exists in the regime where $v^{2}<u^{2}$ [24, 25, 26, 27].
@@
$S=i\int d^{1+1}x|e|\bar{\psi}e^{\mu}_{a}\gamma^{a}\overset{\leftrightarrow}{\partial}_{\mu}\psi(x),$ (8)
where $\overset{\leftrightarrow}{\partial}_{\mu}=\frac{1}{2}(A\partial_{\mu}B-(\partial_{\mu}A)B)$, the coefficients $e^{t}_{0}=1$ and $e^{t}_{1}=1$, the flat space-time gamma matrices $\gamma^{a}=\sigma^{z}\alpha^{a}$, and the conjugate spinor $\bar{\psi}(x)=\psi^{\dagger}(x)\gamma^{0}$ have been defined, and $\mu\in\{t,x\}$. Here, $|e|=\det(e^{a}_{\mu})=\sqrt{|g|}$, where $g$ is the determinant of the covariant metric tensor $g_{\mu\nu}$ [24, 25, 26, 27].
@@
If the couplings $v$ and $u$ are slowly varying position-dependent functions $v(x)$ and $u(x)$, then the line element of Eq. (10) is that of the (1+1)-dimensional Gullstrand-Painlevé metric [39, 40], with the (+)(+\ -) metric signature, which describes the space-time geometry of a (1+1)-dimensional black hole with an event horizon of radius $x_{h}$ located at $v(x_{h})^{2}=u(x_{h})^{2}$ [21, 22]. The Gullstrand-Painlevé metric does not possess a coordinate singularity at $v(x_{h})^{2}=u(x_{h})^{2}$ and, therefore, is valid in both the interior and exterior space-time of the black hole, with the region bounded by $v(x)^{2}\geq u(x)^{2}$ corresponding to the interior [25, 24]. Though the Gullstrand-Painlevé metric is not static (not invariant under the time-reversal transformation $t\mapsto-t$), it is stationary (invariant under the time-translation transformation $t\mapsto t+t_{0}$). As the metric is stationary, it admits a time-like Killing vector $\xi^{\mu}=\delta^{\mu}_{t}$ (equivalently denoted $\partial_{t}$), which allows energy to be well-defined in the black hole’s interior region and conserved along the metric’s geodesics [22, 23, 21] (see Appendix B.).
@@
III Entropy of simulated (1+1)(1+1)D Dirac fermions in black hole background
@@
In this section, we derive analytic expressions for the entropy of the fermion zero-modes in the interior and exterior regions of a $(1+1)$-dimensional black hole. Two assumptions, that the fermions in the black hole’s interior region are in thermal equilibrium with the exterior region and that the temperature of this thermal equilibrium is the Hawking temperature, are key to this derivation. The derivation utilises a non-linear dispersion relation to find a finite density of states of the fermion zero-modes, which, in turn, can be used along with the assumptions of thermal equilibrium to obtain their thermal energy and, hence, thermodynamic entropy. Under this assumption of thermality, we expect that the thermodynamic entropy of these fermion zero-modes is equivalent to the entanglement entropy associated with the bipartitioning of space-time by the event horizon [13, 36]. In Sec. IV, where numerical simulations of the chiral spin-chain’s entanglement entropy are employed, this equivalence will be used to probe the thermality of the black hole.
@@
We expect that, in the low-energy limit where the mean-field theory is valid, the thermodynamic behaviour of the fermions in the black hole’s interior will be determined by the fermionic states near the Fermi energy $E(p)=0$, as all negative energy states will be occupied [22, 23]. The density of states $N(0)$, at zero energy, of the fermion zero-modes enclosed in a sub-region of length $s\leq x_{h}-a_{c}$ of the black hole’s interior will be given by [21, 22]
$N(0)=\frac{N_{F}}{\pi\hbar}\int^{s}_{0}dx\int_{\mathbb{R}}dp\ \delta(\tilde{E}(p)),$ (11)
@@
where $\tilde{E}(p)$ is the dispersion relation of the fermions and $N_{F}$ denotes the integer number of massless fermion species. To prevent the density of states of Eq. (11) from being divergent, we require two conditions to be satisfied. First, the upper limit of $s$ must be restricted to a distance $a_{c}$ from the event horizon $x_{h}$ [43]. This prevents an infinite number of degrees of freedom, from points on either side of the event horizon with a separation of less than $a_{c}$, from contributing to the density of states [43, 44, 33]. Second, the dispersion relation must contain an ultraviolet cut-off, which can be introduced via a non-linear dispersion relation that breaks the effective Lorentz invariance at high momenta or, equivalently, at short distances [22, 23, 21, 28, 30, 31, 32, 29]. This violation of the effective Lorentz invariance allows for the density of states to be finite, as, by introducing a fundamental length scale, it prevents the localisation of an infinite number of degrees of freedom into an arbitrarily small volume [43, 44, 33].
As the effective Lorentz invariance emerges in the low-energy limit of the chiral spin-chain’s continuum limit, we can identify an appropriate non-linear dispersion relation by considering a higher-order expansion of the mean-field theory’s dispersion relation given in Eq. (5). Taylor expanding this dispersion relation about the Fermi point $p_{0}$ to $O(p^{2})$, after renormalising the couplings and setting $u=1$, we obtain
$E(p+p_{0})=-vp\pm\sqrt{p^{2}-\frac{a_{c}^{2}p^{4}}{12}}+O(p^{3})\equiv\tilde{E}(p).$ (12)
@@
We will take the non-linear dispersion relation of the fermions in the black hole space-time to be the positive solution (super-luminal) of that in Eq. (12), which has roots at $p=0$ and $p=\sqrt{\frac{12}{a_{c}^{2}}(1-v^{2})}$. With this non-linear dispersion, and making use of the Dirac delta identity $\delta(f(x))=\sum_{i}\delta(x-x_{i})/|f^{\prime}(x_{i})|$, the density of states of the fermion zero-modes enclosed in the interior region of the black hole can be expressed as
$N(0)=\frac{N_{F}}{\pi\hbar}\int^{s}_{0}dx\left(\frac{1}{|v(x)|}+\frac{1}{2}\Biggl|\frac{1}{\left(1-v(x)\right)}-\frac{1}{\left(1+v(x)\right)}\Biggr|^{-1}\right),$ (13)
@@
which can be solved for a given coupling $v(x)$ to obtain the relevant density of states. Here, we take the coupling $v(x)=\sqrt{x_{h}/x}$, which corresponds to that of a Schwarzschild black hole [21, 22]. For this coupling, the density of states can be solved to yield
@@
Assuming that the fermions in the black hole’s interior are in thermal equilibrium with the exterior space-time, the density of states can be used to find the thermal energy $\mathcal{E}(T)$ and, hence, the thermodynamic entropy $\mathcal{S}(T)$ of the fermions in the interior region of the black hole. The thermal energy of the fermions in the black hole’s interior will be given by [22, 23, 21]
$\mathcal{E}(T)=N(0)\int^{\infty}_{0}d{\tilde{E}}\frac{\tilde{E}}{e^{\tilde{E}/T}+1},$ (15)
which can be solved using the change of variables $u=-\tilde{E}/T$ and the dilogarithmic identity $\text{Li}_{2}(1)=\pi^{2}/6$ [45, 46] to give [22, 23, 21]
@@
Assuming that, in addition to being in thermal equilibrium, the fermions in the black hole’s interior have a temperature equal to the Hawking temperature [5, 6]
@@
$\mathcal{S}(T_{H})=\frac{N_{F}}{12}\ln(\frac{4x_{h}}{a_{c}})-\frac{N_{F}}{9},$ (20)
which is logarithmically divergent with the lattice constant $a_{c}$.
@@
$l_{p}^{2}=\frac{1}{\hbar G_{\text{eff}}}=\frac{N_{F}}{6}\ln(\frac{4x_{h}}{a_{c}}).$ (21)
@@
In addition to studying the fermion zero-mode entropy of the black hole’s interior, we can extend this analysis to derive an expression for the fermion zero-mode entropy of both the black hole’s interior and exterior regions. The method of this derivation is analogous to that of the black hole’s interior, and so, for brevity, we leave the details of this derivation to Appendix D. It involves determining the density of states of Eq. (13) for both the interior region, which is bound by $0,x_{h}-a_{c}]$, and a portion of the exterior space-time, which we take to be bound by $x_{h}+a_{c},s]$, where $s\geq x_{h}+a_{c}$ is the distance from the coordinate origin $x=0$. Here, as before, the lower limit of $s$ has been restricted to a distance $a_{c}$ from the event horizon to prevent degrees of freedom of points on either side of the event horizon with a separation of less than $a_{c}$ from contributing to the density of states and entropy. Once again, assuming that the fermions in the black hole’s interior are in thermal equilibrium with the exterior region and that the temperature of this equilibrium is given by $T_{H}$, the fermions’ thermal energy and thermodynamic entropy can be defined, the latter of which can be solved to yield
$\mathcal{S}(T_{H})=\frac{N_{F}}{12}\ln\left(\frac{4x_{h}}{a_{c}}\right)+\frac{\tilde{N}_{F}}{12}\left(\frac{2}{3}\left(\frac{s}{x_{h}}\right)^{\frac{3}{2}}+2\left(\frac{s}{x_{h}}\right)^{\frac{1}{2}}+\ln\left|\frac{{s}^{\frac{1}{2}}-x_{h}^{\frac{1}{2}}}{{s}^{\frac{1}{2}}+x_{h}^{\frac{1}{2}}}\right|+\ln\left(\frac{4x_{h}}{a_{c}}\right)\right)+\mathcal{S}_{0}$ (22)
in the limit $0<a_{c}\ll x_{h}$, where $\tilde{N}_{F}$ denotes the integer number of massless fermion species in the black hole space-time’s exterior region and $\mathcal{S}_{0}=-(N_{F}-2\tilde{N}_{F})/9$.

### v2

@@ Abstract
We investigate the apparent thermality of Hawking radiation in the semi-classical limit of quantum black holes using the mean-field limit of a chiral spin-chain simulator, which models fermions propagating on a black hole space-time in the continuum. In this free-theory regime, no genuine thermalisation occurs. Nevertheless, we show that a bipartition across the event horizon yields a reduced density matrix whose mode occupations follow an apparent thermal Fermi–Dirac distribution. In contrast, partitions away from the horizon do not exhibit such thermal distributions, reflecting the absence of thermal behaviour. Our results demonstrate that Hawking radiation appears thermal only with respect to horizon bipartitions in free theories, whilst genuine thermal behaviour emerges only in the presence of interactions deep in the black hole interior.
@@ I Introduction
Despite theoretical advancements in the study of quantum black holes, there remains several fundamental open questions, including whether, via black hole formation and evaporation, information of the black hole’s internal content is erased (the black hole information paradox) [29, 1]; why black hole entropy appears to be independent of the number of quantum fields [37, 58, 56]; what quantum degrees of freedom can be attributed to black hole entropy [67, 70]. In an effort to illuminate our understanding of these questions, there have been significant efforts, theoretically and, more recently, experimentally [51, 72, 45, 43, 39, 34, 57, 52], towards realising physical analogues of black holes. Such a system was first proposed by Unruh, who argued that sonic analogues of black holes would experience a process analogous to the Hawking effect [61, 62]. Since then, a myriad of systems have been proposed to simulate not only the physics of black holes but also cosmological inflation, the early universe, and aspects of the standard model [2, 65, 66, 64, 35].
@@
Here, we consider a chiral spin-chain model [33, 32, 22, 18] that, in its continuum limit, due to the emergence of an effective Lorentz invariance in its low-energy limit, describes a theory of non-interacting Dirac fermions in a curved space-time geometry. This Lorentz invariance is naturally broken at the length scales of the spin-chain’s lattice spacing as a consequence of the model’s non-linear dispersion relation. We calculate the density of states of the fermion zero-modes in various sub-regions of the space-time describing a (1+1)-dimensional black hole, which is finite as a consequence of the non-linear dispersion relation [64, 35, 66, 12, 13, 15, 14, 36]. By assuming that the fermions inside the black hole are well described by a thermal state at the Hawking temperature, the density of states can be used to compute the entropy of the fermion zero-modes across different regions of the black hole space-time. This includes the interior, where the entropy diverges logarithmically as the lattice spacing approaches zero [31, 63, 8], as well as the exterior.
@@
In parallel, we employ numerical simulations to investigate the entanglement entropy of the chiral spin-chain following a bipartition of the lattice into two subsystems. Generally, we expect that the entropy of the fermion zero-modes agrees with the entanglement entropy of the chiral spin-chain if the state describing the partitioned system is thermal, as, in this case, the von Neumann entropy is equivalent to the thermodynamic entropy of the subsystem’s microstates [70, 42]. When the system is partitioned at the event horizon, we find that the entanglement entropy is well described by the analytic fermion zero-mode entropy of the black hole’s interior, suggesting that the ground state of the partitioned system is thermal. However, when the system is partitioned further from the horizon, we find that this agreement breaks down, indicating that the subsystem’s ground state appears thermal only when partitioned near the event horizon [11, 6]. We verify this by studying the subsystem’s mode occupation, which reveals a thermal Fermi-Dirac distribution only for horizon partitions. This breakdown of thermality for arbitrary partitions indicates that no genuine thermalisation occurs in the free-theory regime, implying that the simulated black hole’s information is not entirely erased [30].
@@ II Lattice simulator of Dirac fermions in black hole background
In this section, we introduce the chiral spin-chain model [33, 32, 22, 18], for which, due to the emergence of an effective Lorentz invariance at low energies, the continuum limit Hamiltonian is described by that of massless Dirac fermions propagating on a $(1+1)$-dimensional curved space-time geometry. Thus, by simulating the chiral spin-chain Hamiltonian, we can probe the entropy and thermality of fermions on a black hole space-time.
@@
By re-labelling the lattice sites such that they alternate between two sub-lattices, $A$ and $B$, and introducing a two-site unit cell, as illustrated in Fig. 1, the mean-field Hamiltonian of Eq. (2) can, following a Fourier transform, be expressed as [33, 32, 22, 18]
@@
$E(p)=v\sin(a_{c}p)\pm 2u\cos\left(a_{c}p/2\right),$ (5)
which has Fermi points located at momenta $p_{0}=\frac{\pi}{a_{c}}$ and $p_{\pm}=\pm\frac{1}{a_{c}}\arccos(1-\frac{2u^{2}}{v^{2}})$, the latter of which exist only in the regime where $v^{2}>u^{2}$. In Fig. 2, we see that the dispersion relation possesses Dirac cones that are tilted by the coupling $v$, resulting in the emergence of these Fermi points when these cones over-tilt [33, 32, 22, 18].
Figure 1: Schematic of the lattice corresponding to the chiral spin-chain model Hamiltonian of Eq. (1), where the lattice sites alternate between two sub-lattices, $A$ and $B$, and a two-site unit cell has been introduced.
Figure 2: Blue and orange curves show the positive and negative mean-field dispersion relations of Eq. (5), respectively, for fixed $u=1$ and various coupling strengths $v$ (from left to right, v=0, 0.5, 1, 2v=0,\ 0.5,\ 1,\ 2). As the coupling $v$ increases, the Dirac cone about the Fermi point $p_{0}$ tilts. At the critical point $u^{2}=v^{2}$, which corresponds to the event horizon, the cone over-tilts, resulting in the emergence of two further Fermi points $p_{\pm}$.
@@
$S=i\int d^{1+1}x|e|\bar{\psi}(x)e^{\mu}_{a}\gamma^{a}\overset{\leftrightarrow}{\partial}_{\mu}\psi(x),$ (8)
where $A\overset{\leftrightarrow}{\partial}_{\mu}B=\frac{1}{2}(A\partial_{\mu}B-(\partial_{\mu}A)B)$, the coefficients $e^{t}_{0}=1$ and $e^{t}_{1}=1$, the flat space-time gamma matrices $\gamma^{a}=\sigma^{z}\alpha^{a}$, and the conjugate spinor $\bar{\psi}(x)=\psi^{\dagger}(x)\gamma^{0}$ have been defined, $|e|=\det(e^{a}_{\mu})$, and $\mu\in\{t,x\}$ [33, 32, 22, 18].
@@
It can be verified from the lattice spinor’s anti-commutation relations and the zweibein of Eq. (9) that the curved space-time spinor satisfies the necessary anti-commutation relations of Eq. (A.20).
If the couplings $v$ and $u$ are slowly varying position-dependent functions $v(x)$ and $u(x)$, then the line element of Eq. (10) is that of the (1+1)-dimensional Gullstrand-Painlevé metric [26, 46], which describes the space-time geometry of a (1+1)-dimensional black hole when the surface gravity $\kappa$ of the metric is positive, and a white hole when it is negative [66, 64] (see Appendix 1). The Gullstrand-Painlevé metric does not possess a coordinate singularity at the event horizon, which is located at $v(x_{h})^{2}=u(x_{h})^{2}$, and is therefore valid in both the interior and exterior space-time of the black/white hole, with the region bounded by $v(x)^{2}\geq u(x)^{2}$ corresponding to the interior [32, 33]. Though the Gullstrand-Painlevé metric is not static (not invariant under the time-reversal transformation $t\mapsto-t$), it is stationary (invariant under the time-translation transformation $t\mapsto t+t_{0}$). As the metric is stationary, it admits a time-like Killing vector $\xi^{\mu}=\delta^{\mu}_{t}$ (equivalently denoted $\partial_{t}$), which allows energy to be well-defined in the black hole’s interior region and conserved along the metric’s geodesics [64, 35, 66] (see Appendix A).
@@
III Entropy of simulated (1+1)D Dirac fermions in black hole background
@@
In this section, we derive analytic expressions for the entropy of the fermion zero-modes in the interior and exterior regions of a $(1+1)$-dimensional black hole. Two assumptions, that the state describing the fermions in the black hole’s interior region is approximately thermal (that is, well described by a Gibbs state), and that the temperature of this state is the Hawking temperature, are key to this derivation. The derivation utilises a non-linear dispersion relation to find a finite density of states of the fermion zero-modes, which, in turn, can be used along with the assumptions of thermality to obtain their thermal energy and, hence, thermodynamic entropy. Under this assumption of thermality, we expect that the thermodynamic entropy of these fermion zero-modes is equivalent to the entanglement entropy associated with the bipartitioning of space-time by the event horizon [70, 42]. In Sec. IV, where numerical simulations of the chiral spin-chain’s entanglement entropy are employed, this equivalence will be used to probe the thermality of the black hole.
@@
In the mean-field description, the ground state corresponds to a filled Fermi sea with all negative-energy single-particle modes occupied. Fluctuations around this state are described by particles with positive energy and by antiparticles realised as holes created by removing fermions from the occupied negative energy states. Hence, in the low-temperature regime, the thermodynamic behaviour will be determined by excitations near the Fermi energy $E(p)=0$. The density of states $N(0)$, at zero energy, of the fermion zero-modes enclosed in a sub-region of length $s$ of the black hole’s interior will be given by [66, 64]
$N(0)=\frac{N_{F}}{2\pi\hbar}\int^{s}_{a_{c}}dx\int_{\mathbb{R}}dp\ \delta(E(p)),$ (11)
where $E(p)$ is the dispersion relation of the fermions and $N_{F}$ denotes the integer number of massless fermion fields.
@@
As the effective Lorentz invariance emerges in the low-energy limit of the chiral spin-chain’s continuum limit, we can identify the mean-field dispersion relation of Eq. (5) as an appropriate dispersion relation for the fermions in the black hole space-time. Moreover, as the continuum limit is defined microscopically on a lattice, the lattice spacing provides a natural ultraviolet cut-off that prevents the density of states of Eq. (11) from being divergent. Operationally, this enters in two equivalent ways: First, the spatial integral should not be extended to separations smaller than $a_{c}$ from the event horizon $x_{h}$ and the coordinate origin $x=0$; that is, $a_{c}\leq s\leq x_{h}-a_{c}$ [7]. This prevents an infinite number of degrees of freedom, from points on either side of the horizon or the coordinate origin, from contributing to the density of states [7, 9, 31]. Second, as the mean-field dispersion relation is non-linear, it breaks the effective Lorentz invariance at distances comparable to the lattice spacing [64, 35, 66, 12, 15, 14, 36, 13]. This allows for the density of states to be finite, as, by introducing a fundamental length scale, it prevents the localisation of an infinite number of degrees of freedom into an arbitrarily small volume [7, 9, 31].
@@
In the regime corresponding to the black hole’s interior, the mean-field dispersion relation has three zero-modes, all of which are sub-luminal in nature (those of the negative solution of Eq. (5)) when $u$ and $v$ are of the same sign. Thus, taking the sub-luminal mean-field dispersion relation, and making use of the Dirac delta identity $\delta(f(x))=\sum_{i}\delta(x-x_{i})/|f^{\prime}(x_{i})|$, the density of states of Eq. (11) for the fermion zero-modes enclosed in the interior region of the black hole can be expressed as
$N(0)=\frac{N_{F}}{2\pi\hbar}\int^{s}_{a_{c}}dx\left(\frac{1}{|1-v(x)|}+\Biggl|\frac{1}{\left(1-v(x)\right)}-\frac{1}{\left(1+v(x)\right)}\Biggr|\right),$ (12)
@@
where the couplings have been renormalised and we have taken $u=1$, which can be solved for a given coupling $v(x)$ to obtain the relevant density of states. Here, we take the coupling $v(x)=\sqrt{x_{h}/x}$, which corresponds to that of a Schwarzschild black hole [66, 64]. For this coupling, the density of states of Eq. (12) can be solved to yield
@@
Assuming that the reduced state associated with the fermions in the black hole’s interior is well described by a thermal Gibbs state, the density of states can be used to find the thermal energy $\mathcal{E}(T)$ and, hence, the thermodynamic entropy $\mathcal{S}(T)$ of the fermions in the interior region of the black hole. The thermal energy of the fermions in the black hole’s interior will be given by [64, 35, 66]
$\mathcal{E}(T)=N(0)\int^{\infty}_{0}d{E}\frac{E}{e^{E/T}+1},$ (14)
which can be solved using the change of variables u=E/Tu=-E/T and the dilogarithmic identity $\text{Li}_{2}(1)=\pi^{2}/6$ [55, 40] to give [64, 35, 66]
@@
Assuming that, in addition to being described by a thermal state, and as semi-classical intuition suggests, the temperature of this thermal state is equal to the Hawking temperature [27, 28]
@@
$\mathcal{S}(T_{H})=\frac{N_{F}}{6}\ln(\frac{2x_{h}}{a_{c}})+\mathcal{S}_{0},$ (19)
where $S_{0}=N_{F}(2\ln(2)-7)/24$, which is logarithmically divergent with the lattice constant $a_{c}$.
@@
$\frac{1}{l_{p}^{2}}=\frac{1}{\hbar G_{\text{eff}}}=\frac{N_{F}}{3}\ln(\frac{2x_{h}}{a_{c}}).$ (20)
@@
In addition to studying the fermion zero-mode entropy of the black hole’s interior, we can extend this analysis to derive an expression for the fermion zero-mode entropy of both the black hole’s interior and exterior regions. The method of this derivation is analogous to that of the black hole’s interior, and so, for brevity, we leave the details of this derivation to Appendix C. It involves determining the density of states of Eq. (12) for both the interior region, which is bound by $a_{c},x_{h}-a_{c}]$, and a portion of the exterior space-time, which we take to be bound by $x_{h}+a_{c},s]$, where $s\geq x_{h}+a_{c}$ is the distance from the coordinate origin $x=0$. Here, as before, the lower limit of $s$ has been restricted to a distance $a_{c}$ from the event horizon to prevent degrees of freedom of points on either side of the event horizon with a separation of less than $a_{c}$ from contributing to the density of states and entropy. In contrast to before, only the zero-mode corresponding to the Fermi point at $p_{0}$ will contribute to the density of states in the exterior region. Considering this, and once again assuming that the reduced state describing the fermions in the black hole’s interior is suitably described by a thermal Gibbs state with a temperature $T_{H}$, the fermions’ thermal energy and thermodynamic entropy can be defined, the latter of which can be solved to yield
$\mathcal{S}(T_{H})=\frac{N_{F}}{6}\ln\left(\frac{2x_{h}}{a_{c}}\right)+\frac{\tilde{N}_{F}}{12}\left(\frac{1}{2}\left(\frac{s}{x_{h}}\right)+\left(\frac{s}{x_{h}}\right)^{\frac{1}{2}}+\ln\left(\frac{2x_{h}^{\frac{1}{2}}}{a_{c}}\right)+\ln\left|{s}^{\frac{1}{2}}-x_{h}^{\frac{1}{2}}\right|\right)+\mathcal{S}^{\prime}_{0}$ (21)
in the limit $0<a_{c}\ll x_{h}$, where $\tilde{N}_{F}$ denotes the integer number of massless fermionic fields in the black hole space-time’s exterior region and $\mathcal{S}^{\prime}_{0}=\mathcal{S}_{0}-3\tilde{N}_{F}/24$.